OpenStudy (anonymous):

I need serious help with these problems!!!! Assume that the team plays 7 games. If the team is equally likely to win as to lose each game, what is the probability that they win a string of at least 5 games in a row?

OpenStudy (anonymous):

jesus another one. this one easier. only way to get at least 5 game in a row is 1) win all 7 2) win 6 lose 1 3) lose 1 win 6 4)win 1, lose 1, win 1 3) win 5 lose 2 4) lose 2, win 5

OpenStudy (anonymous):

i think this is the only way. and since the probability of winning each game is 1/2, each of these events have probability \[\frac{1}{2^7}\] so if i am not mistaken the answer is \[\frac{6}{2^7}\]

OpenStudy (anonymous):

It did not work

OpenStudy (anonymous):

I thought it would be 1/2^7

Directrix (directrix):

3 (1/2 )^7 + 2 (1/2) ^7 + 1 (1/2) ^7 = 6 (1/2) ^7 = 6/128 = 3/64 = .047 approx